OpenStudy (anonymous):

What is '3.00x + 2.50y < 13.00' in slope-intercept form?

OpenStudy (anonymous):

Slope intercept, (as far as I know) is y=mx + b What you're gonna wanna do is keep the 2.50y on the side it's at right now. And turn it into 2.50y = -3.00x + 13.00. Now you may say, why did the 3.00x turn negative? Well because you moved it to the other side. When you move a number on the different side of the equals sign, it become the opposite of what it already is. (If something is positive, and it moves, it becomes negative)

OpenStudy (rock_mit182):

why don't you see the equation in this form and then put the inequallity; \[y = \frac{ -3.00x }{ 2.50} + 13.00\]

OpenStudy (anonymous):

Thank you guys so much :) I thought the same, but I was double-checking

OpenStudy (rock_mit182):

sorry i forgot :

OpenStudy (rock_mit182):

\[y = \frac{- 3.00x }{ 2.50 } + \frac{ 13.00 }{ 2.50 }\]